OpenStudy (anonymous):

5a - (8-a) = 4a - 10

OpenStudy (anonymous):

I assume you're trying to find the value of a: 5a-(8-a) = 4a -10 5a-8+a=4a-10 get all of the a values on one side and the constants on the other: 5a+a-4a =-10+8 2a=-2 a=-1 you can check the answer by substituting in -1 for a: 5(-1)-(9)) = 4(-1) -10 -5-9=-4-10 -14 = -14